OpenStudy (anonymous):

Calculate the value of c. 7.1 7 14.2 3.5

OpenStudy (anonymous):

can you help me again lol @jim_thompson5910

OpenStudy (anonymous):

\[\frac{ c }{ 5 }=\tan 35,c=5*\tan 35=?\]

OpenStudy (anonymous):

3.5

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

can you help me with one more like this??? @surjithayer

OpenStudy (anonymous):

\[\frac{ c }{ 6 }=\cot 43=\cot \left( 90-47 \right)=\tan 47\] c=6tan 47=?

OpenStudy (anonymous):

6.43

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw